OpenStudy (anonymous):

Your gross weekly salary is $800.00. Your federal income tax deduction is $53. The Social Security tax is 6.2%. The Medicare tax is 1.45% of gross pay. The state tax is 1.5% of gross pay. Each week you have $28.40 deducted for medical insurance and $12.00 for a health savings account. What is your net pay?

OpenStudy (kl0723):

@soccerlove same approach as the one we did before ;)

OpenStudy (kl0723):

first percentage is 6.2 so divide that by 100 and you get 0.062 now multiply 800*0.062 = 49.6 second one is 800*0.0145 = 11.6 third one is 800*0.015= 12 substract 28.40 because of medical insurance and last substract 12 because of health savings

OpenStudy (kl0723):

now lets take 800 and apply this deductions: 800-49.6-11.6-12-28.40-12= $686.4 :)